OpenStudy (anonymous):

Solve for y: y + 8 = 6 y = 3 y = 5 y = -3 y = -5

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

answer should be -2

OpenStudy (anonymous):

yeah thats what i got.. ugh.

myininaya (myininaya):

Hey! Please guide the asker next time. Thanks.

OpenStudy (anonymous):

and i also got -3 for this.... Solve for x: -2(x - 3) = 10 x = -2 x = 2 x = -8 x = 8

OpenStudy (anonymous):

\[y+8=6\] subtract 8 on both sides to isolate y \[y+8-8=6-8\] \[y=?\]

OpenStudy (anonymous):

nevermind on the second question.

OpenStudy (anonymous):

distribute \[-2x+6=10\] in order to solve for x, isolate it by subtracting 6 on both sides \[-2x+6-6=10-6\] \[-2x=4\] we are looking for the value of x, not -2x. so we divide both sides by -2